# Entropy–Smooth Structures on Topological Manifolds

### v2

@@ 4.2. Euclidean space $\mathbb{R}^{n}$
Fix $x_{0}\in\mathbb{R}^{n}$. Let the local information probes be the centred Gaussian measures
@@
To construct a valid entropy coordinate chart satisfying Axiom 3, we use normalized directional coordinates. Define
$f_{i}(y):=\frac{y_{i}-x_{0,i}}{\|y-x_{0}\|},\qquad i=1,\dots,n,$
on a punctured neighborhood of $x_{0}$. These functions are continuous, separate points, and satisfy $f_{i}(x_{0})=0$ in the sense of the limit along rays. For the Gaussian probes, the quadratic entropy coefficients are
$I_{x,\varepsilon}(f_{i},f_{j})=\int_{\mathbb{R}^{n}}\frac{(y_{i}-x_{0,i})(y_{j}-x_{0,j})}{\|y-x_{0}\|^{2}}\,\mu_{x,\varepsilon}(dy),$
which converge as $\varepsilon\to 0$ to the Euclidean inner product on the unit sphere:
$I_{x}(f_{i},f_{j})=\delta_{ij}.$
Thus the information Gram matrix is the identity:
$G_{x}(f_{1},\dots,f_{n})=(I_{x}(f_{i},f_{j}))_{1\leq i,j\leq n}=I_{n},$
which is positive definite. The map
$F(y):=\big(f_{1}(y),\dots,f_{n}(y)\big)$
is a homeomorphism from a punctured neighborhood of $x_{0}$ onto an open subset of $\mathbb{R}^{n}$, and $(f_{1},\dots,f_{n})$ forms an entropy coordinate chart.
@@
Write $S^{1}=\mathbb{R}/2\pi\mathbb{Z}$ with angular coordinate $\theta$. Let $d_{S^{1}}$ denote the shortest–arc distance on $S^{1}$. For $\theta\in S^{1}$ and $\varepsilon>0$ define
$\mu_{\theta,\varepsilon}(d\varphi):=\frac{1}{Z_{\varepsilon}}\exp\!\Big(-\frac{d_{S^{1}}(\theta,\varphi)^{2}}{2\varepsilon}\Big)\,d\varphi,$
where the normalisation constant
$Z_{\varepsilon}:=\int_{S^{1}}\exp\!\Big(-\frac{d_{S^{1}}(\theta,\varphi)^{2}}{2\varepsilon}\Big)\,d\varphi$
ensures that $\mu_{\theta,\varepsilon}$ is a probability measure. By symmetry, $Z_{\varepsilon}$ does not depend on $\theta$, and $\mu_{\theta,\varepsilon}\Rightarrow\delta_{\theta}$ as $\varepsilon\to 0$.
For a bounded continuous function $f:S^{1}\to\mathbb{R}$, we again have
$D\bigl((1+tf)\mu_{\theta,\varepsilon}\,\|\,\mu_{\theta,\varepsilon}\bigr)=\frac{t^{2}}{2}\int_{S^{1}}f(\varphi)^{2}\,\mu_{\theta,\varepsilon}(d\varphi)+o(t^{2})\quad\text{as }t\to 0,$
hence
$I_{\theta,\varepsilon}(f)=\int_{S^{1}}f(\varphi)^{2}\,\mu_{\theta,\varepsilon}(d\varphi),\qquad I_{\theta}(f)=\lim_{\varepsilon\to 0}\int_{S^{1}}f(\varphi)^{2}\,\mu_{\theta,\varepsilon}(d\varphi)=f(\theta)^{2}.$
Fix $\theta_{0}\in S^{1}$, choose a small arc $U\subset S^{1}$ around $\theta_{0}$, and choose a lift of $U$ to an interval in $\mathbb{R}$ so that the local angular coordinate $F:U\to\mathbb{R}$ is given by $F(\varphi)=\varphi$. As in the one–dimensional real case, we can re–center and shift this coordinate so that $F(\theta_{0})\neq 0$; then
$I_{\theta_{0}}(F)=F(\theta_{0})^{2}>0,$
and the $1\times 1$ Gram matrix $G_{\theta_{0}}(F)$ is positive definite. Thus the entropy–smooth structure generated by $\{\mu_{\theta,\varepsilon}\}$ is compatible with the usual smooth structure on $S^{1}$.
@@ 5.1. Existence and Uniqueness
Recall that a function sheaf on a topological space $X$ is an assignment $U\mapsto\mathcal{F}(U)\subseteq\mathbb{R}^{U}$ for every open set $U\subseteq X$, such that $\mathcal{F}$ satisfies the usual sheaf axioms of locality and gluing [moerdijk2003models].
@@ 5.4. Products
Indeed, for mixed terms, write $H_{t,s}(x,y)=1+tF_{i}(x)+sG_{j}(y)$ and note that $Ent_{(x,y),\varepsilon}(t,s)$ factors through the product measure $\mu^{X}_{x,\varepsilon}\otimes\mu^{Y}_{y,\varepsilon}$. Differentiating $H_{t,s}\log H_{t,s}$ twice and evaluating at $(t,s)=(0,0)$ yields the product $F_{i}(x)\,G_{j}(y)$, which vanishes after centering the coordinate functions so that $F_{i}(x)=G_{j}(y)=0$ (by replacing each function with its centered version $f\mapsto f-f(x)$, which does not change any entropy coefficients). Hence $I_{(x,y)}(F_{i},G_{j})=0$.
For completeness we verify the product identities for the quadratic entropy response. Recall that the perturbed measure associated to a bounded continuous function $h$ is
$\nu^{\,h}_{(x,y),\varepsilon,t}=(1+th)\,\mu^{X}_{x,\varepsilon}\otimes\mu^{Y}_{y,\varepsilon},$
so that
$Ent_{(x,y),\varepsilon}(t,h)=D\bigl(\nu^{\,h}_{(x,y),\varepsilon,t}\,\big\|\,\mu^{X}_{x,\varepsilon}\otimes\mu^{Y}_{y,\varepsilon}\bigr)=\int_{X\times Y}(1+th)\log(1+th)\,d\mu^{X}_{x,\varepsilon}\,d\mu^{Y}_{y,\varepsilon}.$
If $h(x,y)=F_{i}(x)$ depends only on the first factor, then
$Ent_{(x,y),\varepsilon}(t,F_{i})=\int_{X}(1+tF_{i}(x))\log(1+tF_{i}(x))\,d\mu^{X}_{x,\varepsilon}\cdot\int_{Y}d\mu^{Y}_{y,\varepsilon},$
so the second factor is $1$ and we obtain
$Ent_{(x,y),\varepsilon}(t,F_{i})=Ent^{X}_{x,\varepsilon}(t,f_{i}),$
hence $I_{(x,y)}(F_{i},F_{i^{\prime}})=I_{x}(f_{i},f_{i^{\prime}})$ after taking the quadratic coefficient and the limit $\varepsilon\to 0$. The same argument with $h(x,y)=G_{j}(y)$ gives $I_{(x,y)}(G_{j},G_{j^{\prime}})=I_{y}(g_{j},g_{j^{\prime}})$.
For the mixed terms, consider a two–parameter perturbation
$h_{t,s}(x,y)=tF_{i}(x)+sG_{j}(y),$
so that
$Ent_{(x,y),\varepsilon}(t,s):=Ent_{(x,y),\varepsilon}(h_{t,s})=\int_{X\times Y}(1+tF_{i}(x)+sG_{j}(y))\log(1+tF_{i}(x)+sG_{j}(y))\,d\mu^{X}_{x,\varepsilon}\,d\mu^{Y}_{y,\varepsilon}.$
Differentiating under the integral sign and evaluating at $(t,s)=(0,0)$, we use $\int F_{i}\,d\mu^{X}_{x,\varepsilon}=0$ and $\int G_{j}\,d\mu^{Y}_{y,\varepsilon}=0$ (centering) to obtain
$\left.\frac{\partial^{2}}{\partial t\,\partial s}\right|_{t=s=0}Ent_{(x,y),\varepsilon}(t,s)=\int_{X\times Y}F_{i}(x)G_{j}(y)\,d\mu^{X}_{x,\varepsilon}\,d\mu^{Y}_{y,\varepsilon}=\Bigl(\int_{X}F_{i}\,d\mu^{X}_{x,\varepsilon}\Bigr)\Bigl(\int_{Y}G_{j}\,d\mu^{Y}_{y,\varepsilon}\Bigr)=0.$
Thus $I_{(x,y)}(F_{i},G_{j})=0$ for all $i,j$, and the claimed product decomposition follows after passing to the limit $\varepsilon\to 0$.
@@ 5.9. Entropy characterisation of diffeomorphisms
By the classical inverse function theorem [lee2013smooth], $F$ is then a local $C^{\infty}$ diffeomorphism. Since $F$ is assumed bijective, the local inverses glue to a global inverse $F^{-1}:Y\to X$. The map $F^{-1}$ is continuous because $F$ is a continuous bijection between Hausdorff second countable manifolds. Moreover, $F^{-1}$ is entropy–smooth by assumption, and hence it is classically smooth by Theorem 5.1(a). Thus $F$ is a global $C^{\infty}$ diffeomorphism.
@@ 6. Conclusion
In this paper we presented an information–theoretic foundation for smooth structures on topological manifolds. The central idea is that differentiability can be characterized entirely by the small–scale entropy response of local information probes. Our axioms formalize this principle through locality, continuity, coordinate non–degeneracy, entropy regularity, and topological compatibility. From these purely information–theoretic primitives, we recovered the full classical $C^{\infty}$ structure: entropy coordinate charts are smooth diffeomorphisms onto their images, entropy transition maps are smooth, and immersions, submersions, products, and diffeomorphisms behave exactly as in ordinary differential geometry. As shown by the structural theorems, the entropy–smooth category is equivalent to the classical smooth category, and the information quadratic response determines the smooth structure uniquely.
It is instructive to compare this with other abstract frameworks for smoothness [sikorski1967differential, frolicher1986linear, moerdijk2003models]. Sikorski differential spaces, $C^{\infty}$–rings, and Frölicher–Kriegl structures begin by postulating the algebra of smooth functions. In contrast, the present framework starts from probabilistic data: the infinitesimal behaviour of the Kullback–Leibler divergence under local perturbations of the information probes. Smoothness is not assumed but emerges from the quadratic entropy form $I_{x}(f)$, which plays the role of a second–order infinitesimal detector for coordinate functions. In this sense, the entropy–smooth viewpoint reinterprets differentiability as an information–theoretic stability property.
Conceptually, this work represents the “zeroth–order” layer of a broader program initiated in our earlier papers on information–theoretic differential geometry. That earlier work demonstrated that diffusion and entropy encode geometric quantities such as isoperimetric profiles, Ricci curvature, curvature bounds, and the infinitesimal Riemannian metric. Here we show that even without any metric or geometric structure, a smooth manifold can be reconstructed entirely from the behaviour of entropy at the smallest scales. Thus entropy–smoothness provides a conceptual bridge linking classical smooth manifolds to diffusion geometry and to the information–theoretic techniques developed previously [sangha2025isoperimetric, sangha2025curvature].
Several directions for future research follow naturally. One is the development of an entropy differential calculus, extending the present axioms to define entropy–gradients, divergence, Hessians, and second–order operators. Another is to apply the entropy–smooth framework in nonsmooth or fractal settings, where the existence of information probes is natural but no classical smooth structure is present. Finally, it is of great interest to examine how entropy–smooth structures interact with the curvature and comparison geometry tools developed in our prior work [sangha2025isoperimetric, sangha2025curvature], and how these perspectives might unify into a single information–theoretic foundation for both smooth and metric geometry.

### v3

@@ 4.2. Euclidean space $\mathbb{R}^{n}$
Fix $x\in\mathbb{R}^{n}$ and take the centred Gaussian probes
@@
Let $f_{i}(y):=y_{i}$ be the standard coordinate functions. Then $\delta_{x,\varepsilon}f_{i}(y)=(y_{i}-x_{i})/\sqrt{\varepsilon}$, and since $\mathrm{Cov}(y_{i}-x_{i},y_{j}-x_{j})=\varepsilon\,\delta_{ij}$ under $\mu_{x,\varepsilon}$ we have
$I_{x,\varepsilon}(f_{i},f_{j})=\mathrm{Cov}_{\mu_{x,\varepsilon}}\bigl(\delta_{x,\varepsilon}f_{i},\delta_{x,\varepsilon}f_{j}\bigr)=\delta_{ij},\qquad I_{x}(f_{i},f_{j})=\delta_{ij}.$
Thus the information Gram matrix is the identity, hence positive definite. Moreover $F=(f_{1},\dots,f_{n})$ is a homeomorphism from any open ball onto an open subset of $\mathbb{R}^{n}$. Therefore $(f_{1},\dots,f_{n})$ is an entropy coordinate chart in the sense of Definition 3.1 (see below), and the entropy–smooth atlas agrees with the classical smooth atlas on $\mathbb{R}^{n}$.
@@
We omit details. In a local angular coordinate, the KL increment response gives the usual one–dimensional chart condition, so the entropy–smooth atlas agrees with the classical smooth structure on $S^{1}$.
@@ 5.1. Existence and Uniqueness
Recall that a function sheaf on a topological space $X$ is an assignment $U\mapsto\mathcal{F}(U)\subseteq\mathbb{R}^{U}$ for every open set $U\subseteq X$, such that $\mathcal{F}$ satisfies the usual sheaf axioms of locality and gluing [8].
@@ 5.4. Products
Indeed, by Definition 2.8 we have
$I_{(x,y),\varepsilon}(h_{1},h_{2})=\mathrm{Cov}_{\mu^{X}_{x,\varepsilon}\otimes\mu^{Y}_{y,\varepsilon}}\bigl(\delta_{(x,y),\varepsilon}h_{1},\,\delta_{(x,y),\varepsilon}h_{2}\bigr).$
If $h_{1}=F_{i}=f_{i}\circ\pi_{X}$ and $h_{2}=G_{j}=g_{j}\circ\pi_{Y}$, then $\delta_{(x,y),\varepsilon}h_{1}$ depends only on the XX–variable and $\delta_{(x,y),\varepsilon}h_{2}$ depends only on the YY–variable. Since the probe factors as a product measure, these two random variables are independent under $\mu^{X}_{x,\varepsilon}\otimes\mu^{Y}_{y,\varepsilon}$, and hence their covariance vanishes:
$I_{(x,y),\varepsilon}(F_{i},G_{j})=0.$
The remaining identities for the pure XX– and YY–terms follow immediately from the same definition, because restriction to one factor simply reduces the covariance to the corresponding covariance on that factor.
Thus mixed entropy coefficients vanish, and the remaining identities for the pure XX– and YY–terms follow immediately from the same formula.
@@ 5.9. Entropy characterisation of diffeomorphisms
By the classical inverse function theorem [7], $F$ is then a local $C^{\infty}$ diffeomorphism. Since $F$ is assumed bijective, the local inverses glue to a global inverse $F^{-1}:Y\to X$. The map $F^{-1}$ is continuous because $F$ is a continuous bijection between Hausdorff second countable manifolds. Moreover, $F^{-1}$ is entropy–smooth by assumption, and hence it is classically smooth by Theorem 5.1(a). Thus $F$ is a global $C^{\infty}$ diffeomorphism.
@@ 6. Conclusion
We have introduced an information–theoretic framework for smooth structures on topological manifolds, based on the small–scale quadratic response of entropy associated with local probability probes. Under a precise set of axioms governing locality, continuity, non–degeneracy, and functional closure, this entropy response identifies admissible coordinate functions, yields entropy coordinate charts, and determines a smooth atlas. We proved that the resulting entropy–smooth structure is equivalent to the classical smooth structure: entropy–smooth functions coincide with smooth functions, entropy coordinate charts are smooth diffeomorphisms onto their images, and the induced category is equivalent to the category of smooth manifolds.
Conceptually, this work shows that smooth structure can be characterized without reference to derivatives, tangent vectors, or coordinate calculus, but instead through the behaviour of information under infinitesimal perturbations. The axioms make explicit the minimal information–theoretic content required to recover smoothness, clarifying which aspects of differentiability are intrinsic and which must be assumed.
The entropy–smooth framework provides a natural bridge between abstract approaches to smooth structure, such as differential spaces and smooth function algebras, and probabilistic or information–theoretic models arising from diffusion and entropy. Future work will extend this framework to entropy–based differential operators and to the information–theoretic reconstruction of Riemannian geometry and curvature.